OpenStudy (anonymous):

find the missing side of the trapezoid. Bass 1 5.4 Km Height 2.2 Km. Area= 7.7Km2

OpenStudy (anonymous):

(b1+b2).5h=a (15.4+x).5(2.2)=7.7